# Virtualizing RAN: Science, Strategy, and Architecture of Software‑Defined Mobile Networks

### v2

@@ Abstract
Virtualising the Radio–Access Network (RAN) is increasingly viewed as an enabler of affordable 5G expansion and a stepping-stone toward AI-native 6G. Most discussions, however, still approach spectrum policy, cloud engineering and organisational practice as separate topics. This paper offers an integrated perspective spanning four pillars—science, technology, business strategy and culture. A comparative U.S. case study illustrates how mid-band contiguity, complemented by selective mmWave capacity layers, can improve both coverage and churn when orchestrated through software-defined carrier aggregation. We derive analytic capacity and latency bounds for Split 7.2$\times$ vRAN/O-RAN deployments, quantify the throughput penalty of end-to-end 256-bit encryption, and show how GPU/FPGA off-load plus digital-twin–driven automation keeps the hybrid-automatic-repeat request (HARQ) round-trip within a 0.5 ms budget. When these technical enablers are embedded in a physics-first delivery roadmap, average vRAN cycle time drops an order of magnitude—even in the presence of cultural head-winds such as “dual-ladder” erosion. Three cybernetic templates—the Clock-Hierarchy Law, Ashby’s Requisite Variety and a delay–cost curve—are then used to explain why silo-constrained automation can amplify, rather than absorb, integration debt. Looking forward, silicon-paced 6G evolution (9–12-month node shrinks, sub-THz joint communication-and-sensing [^1] JCAS = Joint Communication and Sensing., chiplet architectures and optical I/O) calls for a dual-resolution planning grid that couples five-year spectrum physics with six-month “silicon sprints.” The paper closes with balanced, action-oriented recommendations for operators, vendors and researchers on sub-THz fronthaul, AI-native security, energy-proportional accelerators and zero-touch assurance.
Index Terms:
Virtualised RAN, Open RAN, Split 7.2 ×, RAN Intelligent Controller, Digital Twin, Massive MIMO, Carrier Aggregation, Clock-Hierarchy Law, Requisite Variety, Sub-THz, Joint Communication and Sensing, SDN, Mobile Edge Computing, AI-native Security, Zero-Trust, Change Management, 5 G, 6 G.
@@ I-B Scope and Contributions
    AI‑Native Automation Blueprint—Building on open testbeds such as OAIC and X5G, I map how Digital‑Twin pipelines and RIC xApps close the loop between planning, optimisation and self‑healing [14, 15, 16].
@@
    Change‑Management Analysis—I relate technical road‑blocks to organisational mis‑alignments, illustrating lessons from Verizon’s mmWave pivot and T‑Mobile’s mid‑band momentum.
@@ I-C Paper Organization
Positioning within Systems Theory— Having drilled down to FFT-symbol deadlines and carrier–aggregation knapsacks, the remainder of the paper deliberately zooms out to the macro–dynamics that govern how large organisations absorb (or repel) exponential technology. Section VI frames virtualised RAN as an unfolding socio-technical control problem: the Clock-Hierarchy Law shows how strategy outruns execution when internal R&D clocks are allowed to stall; Ashby’s Requisite Variety quantifies the resilience lost when dual-career ladders rust; and the delay–cost curve explains why local automation without a unifying road-map creates compounding integration debt. Culture Eats Strategy for Breakfast: The Nuance of Technology and Infrastructure—adds balance by contrasting “up-or-out” M&A playbooks with pockets of deep technical practice that still thrive inside the same firms, showing that the issue is less which company than how clocks, careers, and capital are synchronised. Taken together, these cybernetic lenses turn the paper from a “how-to” guide for vRAN deployment into a theory-backed playbook for scaling AI-native 6G without replaying 5G’s cultural mis-steps. Readers focused on RF and compute may stop at Section IV; those charged with steering national networks through the next decade should continue into Section VI. Finally, Section VII synthesises the technical and cultural threads and outlines open research directions toward an AI-native, zero-trust, energy-neutral 6G.
@@ II-A Historical Foundations
The organisational DNA of today’s mobile operators can be traced to the pre‑divestiture Bell System monopoly and its subsequent fragmentation into Regional Bell Operating Companies (RBOCs) in 1984. Those incumbents—henceforth Incumbent Local Exchange Carriers (ILECs)—inherited a culture centred on engineering conservation, proprietary hardware, five‑nine availability targets and top‑down capital allocation cycles. While this model delivered unrivalled voice reliability, it also entrenched long replacement cycles and vendor lock‑in that now clash with the software‑driven ethos of virtualized RAN. The deceleration of formal “dual‑career ladders” within many ILECs [^3] A dual ladder permits parallel technical and managerial advancement; its erosion has diverted subject‑matter experts into project‑management roles, eroding in‑house design competence. further weakened internal technology advocacy, tilting power toward cost‑optimised sourcing teams and short‑term revenue goals. Consequently, when 5G emerged, the U.S. incumbents reacted with hardware‑first tactics—Verizon’s mmWave Ultra‑Wideband (UWB) Nationwide Coverage attempt being the canonical example—rather than embracing software abstraction and agile spectrum use.
@@ II-B Spectrum Assets and Auction Dynamics
U.S. spectrum is assigned via ascending‑clock auctions with partial reserve prices; Table I summarises headline events. Notable milestones include Auction 97 (AWS‑3, \$44.9 B, 2015), Auction 107 (C‑band, \$81 B, 2021) and Auction 110 (3.45 GHz, \$22.5 B, 2022) [10]. Each round reshapes operator holdings and debt capacity, dictating where carrier‑aggregation (CA) and Dynamic Spectrum Sharing (DSS) can be deployed profitably. While DSS briefly enabled 5G overlays in 700/850 MHz LTE carriers, operators quietly de‑commissioned most deployments in 2024 after discovering a 20–40% capacity penalty and scheduler complexity [17].
@@ II-C Operator Go‑to‑Market Strategies
Leveraging vast 28/39 GHz holdings, Verizon launched UWB small‑cells to claim multi‑Gb/s marketing headlines. However, limited propagation demanded thousands of additional street‑level poles; CapEx pressure and negative churn forced a pivot toward C‑band, where Verizon secured 160–200 MHz in the top‑46 PEAs [2]. The firm is now re‑optimising its network with Massive‑MIMO C‑band macro cells supplemented by targeted mmWave FWA.
@@
AT&T split its bids between 3.45 GHz and C‑band while leveraging FirstNet (700 MHz) commitments to densify low‑band coverage. Financial leverage from WarnerMedia spin‑offs limits acceleration, making RAN cost optimisation a strategic imperative.
@@ II-D Regulatory and Compliance Head‑winds
New macro or small‑cell sites require National Environmental Policy Act (NEPA) screening and, where applicable, Tribal and historic‑structure consultation. Processing times stretch from weeks (collocation on existing poles) to $>$ 12 months for green‑field towers, directly impacting annual build targets.
@@
Take‑away—Spectrum economics, auction outcomes and non‑technical regulation collectively shape the technical architecture and timing of vRAN/O‑RAN adoption. Operators that translate diverse holdings into software‑defined, AI‑optimised capacity—while navigating compliance at minimum Opex—are positioned to out‑scale hardware‑centric rivals.
@@ III Technical Foundations of Radio Access Virtualization
Virtualization reshapes the radio access network by decoupling time‑critical baseband logic from monolithic, proprietary appliances. Before examining vRAN and O‑RAN abstractions I establish four technical pillars: (i) the classical hardware chain, (ii) spectrum engineering constraints that dictate cell‑site topology, (iii) the evolution from centralised to virtualized RAN splits, and (iv) the service‑management framework introduced by the O‑RAN Alliance.
@@ III-A Canonical RAN Hardware Chain
The DU firmware limits (i) total cell count (18–21 on pre‑vRAN hardware) and (ii) cumulative Absolute Bandwidth (ABW) per DU ($<\!160$ MHz FR1; $>\!400$ MHz FR2 when split across multiple component carriers). When spectrum is highly fragmented, operators face a cellular “knapsack” problem: either deploy an extra DU or disable least‑profitable carriers to stay within software limits.
@@
[…]
@@
[…]
@@
TABLE VII: Hardware‑limited channelisation in band n257 (28 GHz).
@@ III-C From C‑RAN to vRAN and Cloud‑RAN
The scheme scales address space from $10^{4}$ (legacy 7‑digit gNB IDs) to $10^{8}$ yet preserves intuitive grouping: prefixes 4xxx low‑band, 7xxx mid‑band, 9xxx UWB [^4] The 11-digit logical ID is embedded in the 32-bit gNB-ID specified by 3GPP TS 38.413 by coding MMMCCCC as a 20-bit market–vCU prefix and DDDD in the remaining 12 bits—well within the $12^{32}-$ 1 limit.. It is one of many potential addressing schemes, as 3GPP defines a total address space of $2^{32}$ and allows the operator to set the addressing standard. When architecting address spaces for emerging technologies, backward-compability with legacy addressing schemas must be considered to avoid costly network re-integration.
@@ III-D Open RAN Architecture
The O-RAN Alliance formalises an open reference stack, depicted in Fig. 6. At the top sits the Service Management and Orchestration (SMO) framework, a cloud-native life-cycle manager for all O-RAN Network Functions that exposes the O1 interface (NetConf/YANG) toward radio units (RUs) and A1 policy toward the Non-Real-Time RIC. The Non-RT RIC performs AI/ML training, KPI mining and long-term policy optimisation on timescales greater than one second, while the Near-RT RIC exerts closed-loop control over the RAN in the 10 ms–1 s window by running xApps that communicate via the E2 interface with virtualized DUs and CUs (O-DU/O-CU); typical xApps include interference-mitigation agents and carrier-power-allocation (CPA) logic [16]. Beneath these controllers, the Open Fronthaul specification defines the M-Plane and C/U-Plane for a 7.2x functional split, complete with Category-A/B timing caps.
@@
Inter-vendor integration has exposed practical pitfalls. Early field trials encountered a “bring-up spiral” in which (a) an O-DU from Vendor A refused the M-Plane certificate issued by Vendor B’s SMO, (b) the Near-RT RIC mis-aligned time-synchronisation packets and corrupted Slice-ID mappings, and (c) an RU vendor’s Part-15 filing covered only an unencrypted control plane, obliging operators to repeat certification once security patches were applied. Generative-AI-driven digital-twin platforms such as NS-O-RAN and NVIDIA Aerial now model the entire stack, allowing exhaustive regression tests before a single truck rolls.
Key insight—Virtualization does not erase RF physics; it merely moves bottlenecks from rigid hardware into schedulable software pools. Understanding contiguous spectrum math, DU compute ceilings and security/performance trade-offs is prerequisite to any AI-driven, self-optimising 5G/6G RAN.
@@ IV Enabling Technologies: SDN, Edge Compute and AI
Virtualizing the RAN shifts the design centre from bespoke signal‑processing ASICs to cloud‑native software primitives orchestrated at millisecond time‑ scales. Three technology domains make this transition feasible: software‑defined networking to decouple control and forwarding, edge compute hardware to meet HARQ deadlines, and AI/ML tool‑chains to automate optimisation and self‑healing. Robust security engineering then glues the stack together under zero‑trust assumptions.
@@ IV-A Software‑Defined Networking Principles Applied to RAN
[…]
The Near‑RT RIC plays the role of an SDN controller with 10–100 ms decision loops, installing run‑time policies (slice QoS, power caps, PRB masks) via E2AP [3]. Meanwhile the Non‑RT RIC handles long‑horizon optimisation and model training.
@@
Operators expose slice templates—SST=1x,e2eLatency$<$5ms for URLLC,
SST=1e,throughput$>$250Mbps for eMBB—then delegate resource fulfillment to the RIC. A closed‑form Lagrangian can express the slice admission constraint:
@@
[…]
@@
Beyond E2, the fledgling E4 (interface to user‑plane functions) and SDN‑like SBI/N4 in 5G core enable end‑to‑end cross‑layer optimisation, e.g. binding a URLLC radio slice to a URLLC UPF in 20 µs proximity.
@@ IV-C AI/ML Pipelines
OAIC [14] and X5G [15] emulate gNB stacks, fronthaul and core in mininet‑hifi + GPU‑accelerated PHY loops. Training data derives from OTFS / SC‑FDM channel synthesis plus live KPI streaming ($>2$ GB h-1 per site). NS‑O‑RAN extends this with ns‑3 + pyTorch wrappers so that xApps can be prototyped offline before hitting production E2.
@@
Cisco’s evolution of Self‑Optimising Networks (SON) integrates eXplainable AI (XAI) to surface root causes—e.g. “Sector 2 3.5 GHz downlink SINR ↓ because Rx chain B phase noise ↑”—and triggers automated RU swap orders [13].
@@ IV-D Security, Privacy and Trust Management
The O‑RAN spec mandates TLS 1.3 or IPSec ESP‑AES‑256‑GCM across E2, A1, O1. Measurements on the OAIC testbed show a 7–12 % user‑plane throughput hit when L2 segmentation forces crypto per MAC PDU rather than jumbo frames [11]. Off‑loading to AES‑NI/SSE or to SNOW‑V vectorised software reduces the hit to $<$ 3 %.
@@
SNOW‑V FPGA implementations can leak 128‑bit sub‑keys in $<$ 50 power traces if masking is omitted [6]; mitigations include domain‑oriented masking and split‑table S‑boxes [21]. At the protocol layer, bidding‑down attacks strip 256‑bit cipher options during Attach, forcing NEA0 null encryption [5]. RIC policy can monitor S‑MC to detect such downgrades in near‑RT.
@@
Synthesis— SDN abstractions give the RAN a programmable control surface; edge hardware provides deterministic compute; AI/ML pipelines translate KPI torrents into real‑time optimisation; and layered 256‑bit cryptography with hardware masks keeps the open stack defensible. Together they form the operational backbone for AI‑native 6G systems.

### v3

@@ Abstract
[…]
@@ I-B Scope and Contributions
    AI‑Native Automation Blueprint—Building on open testbeds such as OAIC and X5G, I map how Digital‑Twin pipelines and RIC xApps close the loop between planning, optimization and self‑healing [14, 15, 16].
@@
    Change‑Management Analysis—I relate technical road‑blocks to organizational mis‑alignments, illustrating lessons from Verizon’s mmWave pivot and T‑Mobile’s mid‑band momentum.
@@ I-C Paper Organization
Positioning within Systems Theory— Having drilled down to FFT-symbol deadlines and carrier–aggregation knapsacks, the remainder of the paper deliberately zooms out to the macro–dynamics that govern how large organizations absorb (or repel) exponential technology. Section VI frames virtualized RAN as an unfolding socio-technical control problem: the Clock-Hierarchy Law shows how strategy outruns execution when internal R&D clocks are allowed to stall; Ashby’s Requisite Variety quantifies the resilience lost when dual-career ladders rust; and the delay–cost curve explains why local automation without a unifying road-map creates compounding integration debt. Culture Eats Strategy for Breakfast: The Nuance of Technology and Infrastructure—adds balance by contrasting “up-or-out” M&A playbooks with pockets of deep technical practice that still thrive inside the same firms, showing that the issue is less which company than how clocks, careers, and capital are synchronized. Taken together, these cybernetic lenses turn the paper from a “how-to” guide for vRAN deployment into a theory-backed playbook for scaling AI-native 6G without replaying 5G’s cultural mis-steps. Readers focused on RF and compute may stop at Section IV; those charged with steering national networks through the next decade should continue into Section VI. Finally, Section VII synthesises the technical and cultural threads and outlines open research directions toward an AI-native, zero-trust, energy-neutral 6G.
@@ II-A Historical Foundations
The organizational DNA of today’s mobile operators can be traced to the pre‑divestiture Bell System monopoly and its subsequent fragmentation into Regional Bell Operating Companies (RBOCs) in 1984. Those incumbents—henceforth Incumbent Local Exchange Carriers (ILECs)—inherited a culture centred on engineering conservation, proprietary hardware, five‑nine availability targets and top‑down capital allocation cycles. While this model delivered unrivalled voice reliability, it also entrenched long replacement cycles and vendor lock‑in that now clash with the software‑driven ethos of virtualized RAN. The deceleration of formal “dual‑career ladders” within many ILECs [^3] A dual ladder permits parallel technical and managerial advancement; its erosion has diverted subject‑matter experts into project‑management roles, eroding in‑house design competence. further weakened internal technology advocacy, tilting power toward cost‑optimized sourcing teams and short‑term revenue goals. Consequently, when 5G emerged, the U.S. incumbents reacted with hardware‑first tactics—Verizon’s mmWave Ultra‑Wideband (UWB) Nationwide Coverage attempt being the canonical example—rather than embracing software abstraction and agile spectrum use.
@@ II-B Spectrum Assets and Auction Dynamics
U.S. spectrum is assigned via ascending‑clock auctions with partial reserve prices; Table I summarizes headline events. Notable milestones include Auction 97 (AWS‑3, \$44.9 B, 2015), Auction 107 (C‑band, \$81 B, 2021) and Auction 110 (3.45 GHz, \$22.5 B, 2022) [10]. Each round reshapes operator holdings and debt capacity, dictating where carrier‑aggregation (CA) and Dynamic Spectrum Sharing (DSS) can be deployed profitably. While DSS briefly enabled 5G overlays in 700/850 MHz LTE carriers, operators quietly de‑commissioned most deployments in 2024 after discovering a 20–40% capacity penalty and scheduler complexity [17].
@@ II-C Operator Go‑to‑Market Strategies
Leveraging vast 28/39 GHz holdings, Verizon launched UWB small‑cells to claim multi‑Gb/s marketing headlines. However, limited propagation demanded thousands of additional street‑level poles; CapEx pressure and negative churn forced a pivot toward C‑band, where Verizon secured 160–200 MHz in the top‑46 PEAs [2]. The firm is now re‑optimizing its network with Massive‑MIMO C‑band macro cells supplemented by targeted mmWave FWA.
@@
AT&T split its bids between 3.45 GHz and C‑band while leveraging FirstNet (700 MHz) commitments to densify low‑band coverage. Financial leverage from WarnerMedia spin‑offs limits acceleration, making RAN cost optimization a strategic imperative.
@@ II-D Regulatory and Compliance Head‑winds
New macro or small‑cell sites require National Environmental Policy Act (NEPA) screening and, where applicable, Tribal and historic‑structure consultation. Processing times stretch from weeks (collocation on existing poles) to >> 12 months for green‑field towers, directly impacting annual build targets.
@@
Take‑away—Spectrum economics, auction outcomes and non‑technical regulation collectively shape the technical architecture and timing of vRAN/O‑RAN adoption. Operators that translate diverse holdings into software‑defined, AI‑optimized capacity—while navigating compliance at minimum Opex—are positioned to out‑scale hardware‑centric rivals.
@@ III Technical Foundations of Radio Access Virtualization
Virtualization reshapes the radio access network by decoupling time‑critical baseband logic from monolithic, proprietary appliances. Before examining vRAN and O‑RAN abstractions I establish four technical pillars: (i) the classical hardware chain, (ii) spectrum engineering constraints that dictate cell‑site topology, (iii) the evolution from centralized to virtualized RAN splits, and (iv) the service‑management framework introduced by the O‑RAN Alliance.
@@ III-A Canonical RAN Hardware Chain
The DU firmware limits (i) total cell count (18–21 on pre‑vRAN hardware) and (ii) cumulative Absolute Bandwidth (ABW) per DU (<160<\!160 MHz FR1; >400>\!400 MHz FR2 when split across multiple component carriers). When spectrum is highly fragmented, operators face a cellular “knapsack” problem: either deploy an extra DU or disable least‑profitable carriers to stay within software limits.
@@
Table VI lists an AWS‑4 block auctioned as $\texttt{c1},\texttt{c2},\texttt{c3},\texttt{c7}$. If an operator wins c1--c3 it may advertise a single 25 MHz aggregated carrier $\mathsf{C1}$ (2110–2135 MHz) because the spectrum is adjacent. The DU maps three component carriers (CCs) into one NR-CA entity, saving two cell IDs and scheduler contexts.
@@
Assume c2 is lost to a competitor. Now c1 and c3 are separated; the DU must create two non‑aggregated carriers $\mathsf{C1}=10$ MHz and $\mathsf{C2}=5$ MHz, plus an out‑of‑band $\mathsf{C3}=15$ MHz at 2675–2700 MHz. Extra carriers $\rightarrow{}$ extra cells $\rightarrow{}$ higher DU load.
@@
TABLE VII: Hardware‑limited channelization in band n257 (28 GHz).
@@ III-C From C‑RAN to vRAN and Cloud‑RAN
The scheme scales address space from $10^{4}$ (legacy 7‑digit gNB IDs) to $10^{8}$ yet preserves intuitive grouping: prefixes 4xxx low‑band, 7xxx mid‑band, 9xxx UWB [^4] The 11-digit logical ID is embedded in the 32-bit gNB-ID specified by 3GPP TS 38.413 by coding MMMCCCC as a 20-bit market–vCU prefix and DDDD in the remaining 12 bits—well within the $2^{32}-1$ limit.. It is one of many potential addressing schemes, as 3GPP defines a total address space of $2^{32}$ and allows the operator to set the addressing standard. When architecting address spaces for emerging technologies, backward-compability with legacy addressing schemas must be considered to avoid costly network re-integration.
@@ III-D Open RAN Architecture
The O-RAN Alliance formalises an open reference stack, depicted in Fig. 6. At the top sits the Service Management and Orchestration (SMO) framework, a cloud-native life-cycle manager for all O-RAN Network Functions that exposes the O1 interface (NetConf/YANG) toward radio units (RUs) and A1 policy toward the Non-Real-Time RIC. The Non-RT RIC performs AI/ML training, KPI mining and long-term policy optimization on timescales greater than one second, while the Near-RT RIC exerts closed-loop control over the RAN in the 10 ms–1 s window by running xApps that communicate via the E2 interface with virtualized DUs and CUs (O-DU/O-CU); typical xApps include interference-mitigation agents and carrier-power-allocation (CPA) logic [16]. Beneath these controllers, the Open Fronthaul specification defines the M-Plane and C/U-Plane for a 7.2x functional split, complete with Category-A/B timing caps.
@@
Inter-vendor integration has exposed practical pitfalls. Early field trials encountered a “bring-up spiral” in which (a) an O-DU from Vendor A refused the M-Plane certificate issued by Vendor B’s SMO, (b) the Near-RT RIC mis-aligned time-synchronization packets and corrupted Slice-ID mappings, and (c) an RU vendor’s Part-15 filing covered only an unencrypted control plane, obliging operators to repeat certification once security patches were applied. Generative-AI-driven digital-twin platforms such as NS-O-RAN and NVIDIA Aerial now model the entire stack, allowing exhaustive regression tests before a single truck rolls.
Key insight—Virtualization does not erase RF physics; it merely moves bottlenecks from rigid hardware into schedulable software pools. Understanding contiguous spectrum math, DU compute ceilings and security/performance trade-offs is prerequisite to any AI-driven, self-optimizing 5G/6G RAN.
@@ IV Enabling Technologies: SDN, Edge Compute and AI
Virtualizing the RAN shifts the design centre from bespoke signal‑processing ASICs to cloud‑native software primitives orchestrated at millisecond time‑ scales. Three technology domains make this transition feasible: software‑defined networking to decouple control and forwarding, edge compute hardware to meet HARQ deadlines, and AI/ML tool‑chains to automate optimization and self‑healing. Robust security engineering then glues the stack together under zero‑trust assumptions.
@@ IV-A Software‑Defined Networking Principles Applied to RAN
$\text{State}\;(KPIs)\;\;\;\xrightarrow{\;\text{A1}/E2\;}\;\;\;\text{Near-/Non‑RT Controllers}\;\;\;\xrightarrow{\;\text{REST/gRPC}\;}\;\;\;\text{RAN\; Node\; Functions}.$
The Near‑RT RIC plays the role of an SDN controller with 10–100 ms decision loops, installing run‑time policies (slice QoS, power caps, PRB masks) via E2AP [3]. Meanwhile the Non‑RT RIC handles long‑horizon optimization and model training.
@@
Operators expose slice templates—SST=1x,e2eLatency<<5ms for URLLC,
SST=1e,throughput>>250Mbps for eMBB—then delegate resource fulfillment to the RIC. A closed‑form Lagrangian can express the slice admission constraint:
@@
where $\mathbf{p}$ is per‑UE power and $R_{u}$ spectral efficiency. Dual ascent runs inside an xApp; the RIC only pushes resulting P_MAX values to vDUs—mirroring SDN “intent $\rightarrow{}$ compiled flow” pipelines.
@@
Beyond E2, the fledgling E4 (interface to user‑plane functions) and SDN‑like SBI/N4 in 5G core enable end‑to‑end cross‑layer optimization, e.g. binding a URLLC radio slice to a URLLC UPF in 20 µs proximity.
@@ IV-C AI/ML Pipelines
OAIC [14] and X5G [15] emulate gNB stacks, fronthaul and core in mininet‑hifi + GPU‑accelerated PHY loops. Training data derives from OTFS / SC‑FDM channel synthesis plus live KPI streaming (>2>2 GB h-1 per site). NS‑O‑RAN extends this with ns‑3 + pyTorch wrappers so that xApps can be prototyped offline before hitting production E2.
@@
Cisco’s evolution of Self‑Optimizing Networks (SON) integrates eXplainable AI (XAI) to surface root causes—e.g. “Sector 2 3.5 GHz downlink SINR ↓ because Rx chain B phase noise ↑”—and triggers automated RU swap orders [13].
@@ IV-D Security, Privacy and Trust Management
The O‑RAN spec mandates TLS 1.3 or IPSec ESP‑AES‑256‑GCM across E2, A1, O1. Measurements on the OAIC testbed show a 7–12 % user‑plane throughput hit when L2 segmentation forces crypto per MAC PDU rather than jumbo frames [11]. Off‑loading to AES‑NI/SSE or to SNOW‑V vectorized software reduces the hit to << 3 %.
@@
SNOW‑V FPGA implementations can leak 128‑bit sub‑keys in << 50 power traces if masking is omitted [6]; mitigations include domain‑oriented masking and split‑table S‑boxes [21]. At the protocol layer, bidding‑down attacks strip 256‑bit cipher options during Attach, forcing NEA0 null encryption [5]. RIC policy can monitor S‑MC to detect such downgrades in near‑RT.
@@
Synthesis— SDN abstractions give the RAN a programmable control surface; edge hardware provides deterministic compute; AI/ML pipelines translate KPI torrents into real‑time optimization; and layered 256‑bit cryptography with hardware masks keeps the open stack defensible. Together they form the operational backbone for AI‑native 6G systems.